OpenStudy (anonymous):

How to evaluate the integral of ∫[x- (1/2x)]^2 dx

OpenStudy (nincompoop):

you can expand

OpenStudy (nincompoop):

then +/- of integral

OpenStudy (anonymous):

is your question written like this one. \[\huge ∫[x- \frac{1}{2x}]^2\] @justine97

OpenStudy (nincompoop):

most likely

OpenStudy (anonymous):

Why did you put evaluate if you have no bounds? o.O that's just an indefinite integral, if you're asking for that just put what is the antiderivative of.. xD

OpenStudy (nincompoop):

you can evaluate without bounds you just put + C

OpenStudy (nincompoop):

he just forgot to put dx. honest mistake

OpenStudy (nincompoop):

\[\int\limits(x^2-1+\frac{ 1 }{ 4x^2 })dx\]

OpenStudy (anonymous):

\[\huge ∫[x- \frac{1}{2x}]^2 dx =\huge ∫[x^2+ (\frac{1}{2x})^2 -2 \times x \times \frac{1}{2x}]\] dx \[ =∫[x^2+ \frac{1}{4x^2} -1] dx = ∫x^2 dx + \int\limits \frac{1}{4x^2} dx -\int\limits 1dx\] \[= ∫x^2 dx + \frac{1}{4} \int\limits x^{-2} dx -\int\limits 1dx\] \[= \frac{x^3}{3} + \frac{1}{4} (\frac{x^{-1}}{-1}) -x +c\] \[= \frac{x^3}{3} - (\frac{x^{-1}}{4}) -x +c\] \[= \frac{x^3}{3} - \frac{1}{4x} -x +c\] @justine97

OpenStudy (nincompoop):

\[\frac{ x^3 }{ 3 }-x+\frac{ 1 }{ 4 }\frac{ x^{-1} }{ -1 } + C\] just simplify

OpenStudy (anonymous):

That guy actually took the time to write all of it in LaTeX xD

OpenStudy (anonymous):

@AshleyAlley<3 thank you

OpenStudy (nincompoop):

ya I wouldn't do that unless it's a complicated demonstration

OpenStudy (nincompoop):

but our student left us without a word or hint of life